(1) find the odds for an event E, when Pr(E) =1/4
(2) find the odds against an event E, when Pr(E)=5/6
(3) find the probability of an event when the odds for the event are 6:4
Help me to understand, please

Answer by ewatrrr(24785)

(1) find the odds for an event E, when Pr(E) =1/4 Odds for: 1:3 4-1+=+3
(2) find the odds against an event E,
when Pr(E)=5/6 P(notE) = 1/6, Odds against: 1:5 6-1+=+5
(3) find the probability of an event,
when the odds for the event are 6:4 P = 6/10 10-6+=+4